1 mol of carbon-12 contains 6.022 x 1023 carbon-12 atoms with a mass of 12 g.
A. Calculate the mass of one carbon-12 atom. Express your answer using scientific notation.
B. Calculate the number of carbon-12 atoms in 20.0 g of carbon-12.
C. Calculate the number of moles of carbon-12 atoms in 20.0 g of carbon-12.

Answer by Boreal(15235)

12 gm/6.022*10^23 = x gm/1
6.022*10^23atoms=12 gm
1 atom=12/6.022 x!0^23=
1.993*10^-23 gm/C atom
===================
20/12*{6.022 x !0^23)=10.037 x 10^23 atoms=1.0037 x 10^24 atoms
==================
That is 5/3 moles.